# TeaMs-RL: Teaching LLMs to Generate Better Instruction Datasets via Reinforcement Learning

### v3

@@ Abstract
TeaMs-RL: Teaching LLMs to Generate Better Instruction Datasets via Reinforcement Learning
Shangding Gu1, Alois Knoll1, Ming Jin2thanks: Corresponding to jinming@vt.edu, shangding.gu@tum.de 1 Technical University of Munich, Munich, Germany 2 Virginia Tech, Virginia, USA
@@ 1 Introduction
Figure 1: Comparative overview of LLM alignment techniques. Current methods (red shaded region) typically involve a two-phase process, starting with Supervised Fine-Tuning (SFT) of a pre-aligned LLM using a dataset of human-crafted instructions and corresponding responses (often sourced from an expert LLM like ChatGPT), leading to a post-SFT LLM. This is then fine-tuned using RLHF, where human feedback on preferences is incorporated, resulting in a post-RLHF LLM. In contrast, our TeaMs-RL method (green shaded region) employs a single-phase SFT approach, initially utilizing RL for teaching expert LLMs to generate high-quality instructions. We train an RL policy (instructor LLM) to create diverse instructions (with the diversity evaluated by a reviewer LLM as a reward signal). Once trained, the instructor LLM produces a set of actions to teach an expert LLM to generate high-quality instructions, and the instructions are leveraged to query the expert LLM to form the SFT instruction dataset. This approach capitalizes on the strengths of RL to enhance the complexity of instructions and consequently the value of responses from the expert LLM.
In this research, we propose a novel method to improve instruction quality with the principles of Reinforcement Learning (RL)  (Sutton and Barto, 2018), hence enhancing LLMs’ ability to comprehend and effectively execute intricate instructions without human involvement, while also boosting model privacy protection. Our method, outlined in the green shaded region of Fig. 1, begins with training an instructor LLM (RL policy) for teaching an expert to generate diverse, complex instructions. These are then used to elicit expert LLM responses, forming a diversity-enhanced instruction dataset. The final stage of supervised fine-tuning (SFT) of a pre-alignment LLM (e.g., Llama-1) with this dataset enhances its complex task processing capabilities. Notably, our study shows that the thoughtful framing of instructions (using RL) is equally, if not more, crucial than generating responses from external sources such as expert LLMs or human feedback.
@@ 2 Related Work and Contribution
Conversely, the self-instruct approach (Wang et al., 2022) charts a different course by tapping into the potential of LLMs to autonomously generate both instructions and their responses, facilitated by the provision of predefined seeds. While this strategy alleviates the dependence on human effort, it might not consistently capture the breadth and depth of diverse instructions and responses typically achieved with human annotators. Building on the pioneering self-instruct methodology, Xu et al. (2024) introduced WizardLM, an evolutionary instruction approach. In WizardLM, initial instructions drawn from the “Evol Alpaca” dataset undergo adaptation through the amalgamation of command instructions and expert LLMs such as ChatGPT. Owing to its commendable performance in both formulating and adhering to intricate instructions, WizardLM has garnered considerable attention. However, its reliance on random sampling of command instructions could fall short in improving the breadth and richness of instructions fed to the LLMs. Furthermore, the heavy dependence on external expert models—commonplace in prevailing self-instruct methods—poses concerns not just economically, but also in terms of environmental impact.
Departing from these precedents, our methodology uniquely uses RL—to help generate foundational fine-tuning instruction data rather than for post-tuning refinements (such as RLHF (Stiennon et al., 2020; Ouyang et al., 2022)). Our key technical contribution is the formulation of a Markov Decision Process (MDP) to train an instructor LLM as an RL policy, adeptly tailored for contextualized instruction manipulations, thereby enabling alignment through SFT on small, high-quality datasets. In contrast to WizardLM’s approach of treating text manipulations as ordinal selections, our MDP framework encodes them in a continuous action space, enabling a more refined differentiation of the intricate effects of textual manipulations (carried out by the instructor LLM) on the quality of instructions. Also, unlike the intricate task of crafting well-structured tree instructions required by Tree-Instruct Zhao et al. (2023b), our MDP can be resolved using TRPO Schulman et al. (2015) (with other common methods being applicable). This approach alleviates the combinatorial complexity inherent in sequential instruction actions, allowing for iterative policy refinement (with TRPO offering the added benefit of guaranteed monotonic progression). Empirically, we demonstrate the efficacy of SFT alignment on a small but high-quality dataset across benchmarks such as ARC and HellaSwag with foundational pre-aligned models such as Llama-1 and Llama-2 (refer to Section 4 for details). Additionally, comparative experiments on model privacy attacks reveal that our method markedly enhances model privacy protection (Section 4.4).
@@ 3 Method
We first train an instructor LLM policy (Sec. 3.1) based on a continuous action space encoding (Sec. 3.1.1) and diversity rules as a reward function (Sec. 3.1.2), a strategy designed to foster the generation of high-quality instructions with an expert LLM such as ChatGPT. Subsequently, using this policy and the expert LLM, we create an instruction-response dataset to fine-tune a pre-aligned LLM. Importantly, our enriched dataset allows bypassing the usual RLHF phase, resulting in a post-SFT LLM that is already proficient in responding to complex instructions. Also, the instruction policy is transferable for aligning various foundation LLMs such as Llama-1-chat-7b and Llama-2-chat-7b.
@@ 3.1 Training the Instructor LLM
To further illustrate how to leverage RL to train the instructor, we first model the RL process with a Markov Decision Process (MDP). An MDP is present as $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, the transition probability function is $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$, the reward function is denoted as $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ and $\gamma$ denotes the discount factor. In an MDP, we aim to maximize the expected cumulative reward values, which are defined as $Q^{\pi}(s,a)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r\left(s_{t},a_{t}\right)\bigg{|}\pi,s_{0}=s,a_{0}=a\right]$ and $V^{\pi}(s)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r\left(s_{t},a_{t}\right)\bigg{|}\pi,s_{0}=s\right]$ for state-action pairs and states, respectively. Moreover, we have the advantage function $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$ Thus, the learning problem is formulated as Equation (1), where $\rho$ denotes an initial state distribution.
@@
In our framework, reward values are assigned by reviewer LLMs, which assess the diversity of actions within a trajectory. A trajectory with highly diverse actions receives a higher reward. The state $s_{t}\in\mathcal{S}$ on a trajectory $\mathcal{T}$ represents the generative instructions. Action space $\mathcal{A}$ is designed in Sec. 3.1.1, and specifical reward settings are provided in Sec. 3.1.2. On the basis of these settings, we can employ a model-free reinforcement learning algorithm (e.g., TRPO Schulman et al. (2015)) to train the instructor LLM.
@@ 3.1.1 Action Set
Inspired by WizardLM (Xu et al., 2024), we leverage several actions to generate complex instructions, including “breadth action”, “add constraint”, “deepening”, “concretizing”, “increase reasoning steps”, “complicate input” (Xu et al., 2024). However, distinct from WizardLM, we map this discrete set into a continuous space by using a language encoding to represent each action, which is then used for a policy to generate instructions. This enables inherently capturing contextual nuances for direct comparison between actions via their Q values. The details of each action are in Appendix B.
@@ 3.1.2 Reward Settings
In this section, consider a judicial prompt $g$. If $g$ is characterized as “equal”, the corresponding reward $r$ is designated as 0. Otherwise, $r$ is set to 1. It is noteworthy that the judgment is determined by the reviewer LLM. For instance, when congruence exists between the initial and subsequent instructions, the reward $r$ is 0. In contrast, a lack of congruence yields a reward value of 1.
For example, a judicial prompt = “““ Here are two Instructions to ChatGPT AI, do you think they are equal to each other, which meet the following requirements: 1. They have same constraints and requirments. 2. They have same depth and breadth of the inquiry. The First Instruction: “““ +instruction + ”””. The Second Instruction: “““ + state + ”””. Your Judgement (Must Just only answer: Equal or Not Equal. No need to explain the reason; ‘Equal’ and ‘Not Equal’ are not allowed to appear simultaneously.): ”””.
@@ 3.2 Fine-tune a Pre-Aligned LLM
As delineated in Fig. 1, an approach is adopted to optimize cost-efficiency considerations during the phase dedicated to fine-tuning a pre-aligned LLM. Specifically, we use the initial instructions to the Alpaca dataset.
The fine-tuning process is executed as follows: We first employ the Alpaca dataset [^4] https://huggingface.co/datasets/tatsu-lab/alpaca as the initial set of instructions and input the initial instructions into the expert LLM like ChatGPT. Next, our trained instruction generation policy, the instructor LLM, will teach the expert LLM to generate high-quality instructions using the initial instructions. Then, the expert LLM produces responses to these high-quality instructions. With the instructions and responses in hand, we fine-tune the pre-aligned LLMs like Llama-1-chat-7b and Llama-2-chat-7b via a supervised fine-tuning way (SFT)—a cost-effective way to enhance LLM capabilities. Finally, we can have the post-SFT LLMs, such as TeaMs-RL-1-7b.
@@ 3.3 Practical Algorithms
The practical training process is outlined in Algorithm 1, which describes the training pipeline. A more detailed algorithm can be found in Algorithm 2, provided in Appendix A. In our training process, in order to teach expert LLMs how to generate high-quality instructions, we leverage the RL policy we have trained to prompt expert LLMs to generate instructions step by step. Building on the training process, we fine-tuned an LLM called “TeaMs-RL”. Within this, we have the flexibility to choose RL methods for policy training, specifically Trust Region Policy Optimization (TRPO) (Schulman et al., 2015) (TRPO is not the only choice), though it has longer training times versus alternatives like Proximal Policy Optimization (PPO). We select TRPO for its rigorous advantage function handling and theoretical guarantees of monotonic improvement.
@@ 4.1 Enhancement of Instruction Diversity
Figure 2: Compare our method with WizardLM in terms of data diversity.
We deploy a policy that is designed to orchestrate a trajectory consisting of six distinct instruction actions, in accordance with the training process to facilitate the generation of instructions by expert LLMs. Note, the expert LLMs utilized in our study, as well as those used by WizardLM, are ChatGPT-3.5 and ChatGPT-4. Both models were accessed in 2023, ensuring that we employed the same versions of the ChatGPT-3.5 and ChatGPT-4 models to generate instructions and responses.. A noteworthy aspect of our investigation involves a comparative assessment of the data quality resulting from the utilization of our policy, in contrast to the approach adopted by WizardLM (Xu et al., 2024), which relies upon random sampling for data generation by querying expert LLMs. It bears emphasizing that our instruction actions are randomly initialized, but as training progresses, our policy iteratively learns to enable the instructor LLM to produce increasingly complex and diverse instructions.
As shown in Fig. 2, our method reliably enhances the diversity score of the instruction set. [^5] In designing instruction actions, the “breadth action” is a single action regarding breath thoughts. Thus, we insert the action into the middle of our trajectory to enhance breadth instructions after training a policy. The main computation overhead is learning the instruction policy on the relatively small WizardLM-13b in less than 1 hour on 2 NVIDIA RTX A6000 GPUs (with 896 total queries). This results in a transferable policy that reduces alignment cost across models compared to RLHF’s per-model RL. Offloading to policy learning provides an instruction set for joint tuning and alignment—a substantial benefit over tuning-only data usage.
4.2 Teach LLMs to Generate Instructions via Trained Policy
To teach expert LLMs in the generation of high-quality instructions, we employ an RL policy that we have trained. This policy prompts the expert LLMs, enabling them to produce high-quality instructions incrementally. This process is achieved by synergizing the initial instructions and ensuring the output enhances each step. This approach facilitates the expert LLMs in refining their instructional outputs, harnessing the RL policy for optimal effectiveness in instruction generation. The related cases are in subsections 4.2.1 and 4.2.2. It is noteworthy that these initial instructions are culled from a well-established and widely-recognized dataset, namely the Alpaca dataset [^6] https://github.com/tatsu-lab/alpaca_eval.git. The experiment settings are available in Appendix I.
4.2.1 Teaching ChatGPT-4 to Generate Imaginative Instructions
For more cases, please see Appendix C.
4.2.2 Teaching ChatGPT-3.5 to Generate Imaginative Instructions
An initial instruction: “Find a unique way to use a rubber band.”
For more cases, please see Appendix D.
@@ 4.3 Compare with WizardLM-7b on ARC and HellaSwag Benchmakrs
To comprehensively examine the effectiveness of our method, we carry out experiments on popular benchmarks: (1) AI2 Reasoning Challenge (ARC) benchmark (Clark et al., 2018): The benchmark introduces a fresh question set, text corpus, and baselines, all strategically curated to foster and propel AI research in the realm of advanced question answering, setting a significantly higher bar for knowledge and reasoning capabilities compared to previous challenges. (2) HellaSwag benchmark (Zellers et al., 2019): The benchmark introduces a challenging dataset, revealing that even state-of-the-art models struggle with commonsense inference, as evidenced by the significant performance gap between humans ($95\%$ accuracy) and models ($48\%$), achieved through adversarial filtering, a robust data collection paradigm that selects adversarial machine-generated wrong answers by scaling up the length and complexity of dataset examples to a ‘Goldilocks’ zone where the text generated is absurd to humans yet often misclassified by models.
We trained a llama-1-7b model denoted “TeaMs-RL-7b-v1.1" with our dataset of 17,878 instruction-response pairs, the training time is about 2 hours on 4 NVIDIA RTX A6000 GPUs. As shown in Figures 3 and 4, our method shows superior performance over WizardLM-7b models in the same experimental settings. It is crucial to highlight an important aspect of our methodology in relation to data utilization. The dataset employed for training our model is approximately one-fourteenth the size of the dataset utilized by WizardLM, as illustrated in Figure 4 (a). Furthermore, Fig. 4 (b) highlights the discernible difference in the query count posed to ChatGPT models between our method and WizardLM, with the latter requesting ChatGPT for responses more than seventeen times than TeaMs-RL. This marked contrast underscores the cost-effectiveness of our data generation approach, which mitigates the expenses associated with dataset acquisition and generates high-quality data. It highlights that our method is a more economically viable and sustainable strategy for training LLMs. For more experimental results, please see Appendix E. Note that the ChatGPT models we used to query data show performance equivalent to the ChatGPT models used in the baselines.
Figure 3: Compare our method with WizardLM 7B on LM-Eval Benchmarks (the higher the value, the better the method’s performance).
Figure 4: Compare with WizardLM 7B on dataset size used for training LLMs (a) and querying count of advanced LLMs (b) (the lower the value, the better the method’s performance).
@@
4.4 Experiments of Model Privacy Attack

### v4

@@ 1 Introduction
Figure 1: Comparative overview of LLM alignment techniques. Current methods (red shaded region) typically involve a two-phase process, starting with Supervised Fine-Tuning (SFT) of a pre-aligned LLM using a dataset of human-crafted instructions and corresponding responses (often sourced from an expert LLM like ChatGPT), leading to a post-SFT LLM. This is then fine-tuned using RLHF, where human feedback on preferences is incorporated, resulting in a post-RLHF LLM. In contrast, our TeaMs-RL method (green shaded region) employs a single-phase SFT approach, initially utilizing RL for teaching expert LLMs to generate high-quality instructions. We train an RL policy (instructor LLM) to create diverse instructions (with the diversity evaluated by a reviewer LLM as a reward signal). Once trained, the instructor LLM produces a set of actions to teach an expert LLM to generate high-quality instructions, and the instructions are leveraged to query the expert LLM to form the SFT instruction dataset. This approach capitalizes on the strengths of RL to enhance the complexity of instructions and consequently the value of responses from the expert LLM. Note that the expert LLMs are not involved in the training of the RL policy; we only use the instructor and the reviewer LLM for training the RL policy. The expert LLM is used to generate instructions and corresponding responses under the guidance of the trained RL policy.
In this research, we propose a novel method to improve instruction quality with the principles of Reinforcement Learning (RL)  (Sutton & Barto, 2018), hence enhancing LLMs’ ability to comprehend and effectively execute intricate instructions without human involvement, while also boosting model privacy protection. Our method, outlined in the green shaded region of Fig. 1, begins with training an instructor LLM (RL policy) for teaching an expert to generate diverse, complex instructions. These are then used to elicit expert LLM responses, forming a diversity-enhanced instruction dataset. The final stage of supervised fine-tuning (SFT) of a pre-alignment LLM (e.g., Llama-1) with this dataset enhances its complex task processing capabilities. Notably, our study shows that the thoughtful framing of instructions (using RL) is equally, if not more, crucial than generating responses from external sources such as expert LLMs or human feedback. Moreover, to clarify our methodology and facilitate replication, we provide a detailed example illustrating the inputs and outputs at each stage of our process, similar to Figure 1, see Appendix A.
@@ 2 Related Work and Contribution
Conversely, the self-instruct approach (Wang et al., 2022) charts a different course by tapping into the potential of LLMs to autonomously generate both instructions and their responses, facilitated by the provision of predefined seeds. While this strategy alleviates the dependence on human effort, it might not consistently capture the breadth and depth of diverse instructions and responses typically achieved with human annotators. Building on the pioneering self-instruct methodology, Xu et al. (2024) introduced WizardLM, an evolutionary instruction approach. In WizardLM, initial instructions drawn from the “Evol Alpaca” dataset are refined through the integration of command instructions and expert LLMs such as ChatGPT. Owing to its commendable performance in both formulating and adhering to intricate instructions, WizardLM has garnered considerable attention. However, its reliance on random sampling of command instructions could fall short in improving the breadth and richness of instructions fed to the LLMs. Furthermore, the heavy dependence on external expert models—commonplace in prevailing self-instruct methods—poses concerns not just economically, but also in terms of environmental impact.
Departing from these precedents, our methodology uniquely uses RL—to help generate foundational fine-tuning instruction data rather than for post-tuning refinements (such as RLHF (Stiennon et al., 2020; Ouyang et al., 2022)). Our key technical contribution is the formulation of a Markov Decision Process (MDP) to train an instructor LLM as an RL policy, adeptly tailored for contextualized instruction manipulations, thereby enabling alignment through SFT on small, high-quality datasets. In contrast to WizardLM’s approach of treating text manipulations as ordinal selections, our MDP framework encodes them in a continuous action space, enabling a more refined differentiation of the intricate effects of textual manipulations (carried out by the instructor LLM) on the quality of instructions. Also, unlike the intricate task of crafting well-structured tree instructions required by Tree-Instruct (Zhao et al., 2023b), our MDP can be resolved using TRPO (Schulman et al., 2015) (with other common methods being applicable). This approach alleviates the combinatorial complexity inherent in sequential instruction actions, allowing for iterative policy refinement (with TRPO offering the added benefit of guaranteed monotonic progression). Empirically, we demonstrate the efficacy of SFT alignment on a small but high-quality dataset across benchmarks such as ARC and HellaSwag with foundational pre-aligned models such as Llama-1 and Llama-2 (refer to Section 4 for details). Additionally, comparative experiments on model privacy attacks reveal that our method markedly enhances model privacy protection (Section 4.3).
@@ 3 Method
We first train an instructor LLM policy (Sec. 3.1) based on a continuous action space encoding (Sec. 3.1.1) and diversity rules as a reward function (Sec. 3.1.2), a strategy designed to foster the generation of high-quality instructions with an expert LLM such as ChatGPT. Subsequently, using this policy and the expert LLM, we create an instruction-response dataset to fine-tune a pre-aligned LLM. Importantly, our enriched dataset allows bypassing the usual RLHF phase, resulting in a post-SFT LLM that is already proficient in responding to complex instructions. Also, the instruction policy is transferable for aligning various foundation LLMs such as Llama-1-chat-7b and Llama-2-chat-7b. Note that our method has been primarily tested on Llama models. There may be performance variations when applying our method to other models. This includes considerations regarding differences in model architecture and computational requirements, which could impact the effectiveness of our approach when transferred to models beyond Llama.
Figure 2: RL policy search for LLM instruction generation. A denotes actions (prompts), S denotes states (instructions), and the green line indicates the optimal instruction generation achieved during the policy search.
@@ 3.1 Training the Instructor LLM
To further illustrate how to leverage RL to train the instructor, we first model the RL process with a Markov Decision Process (MDP). An intuitive example is shown in Figure 2, it illustrates RL policy search for LLM instruction generation, where A represents actions (prompts), S represents states (instructions), and the green path denotes the optimal instruction generation. An MDP is present as $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, the transition probability function is $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$, the reward function is denoted as $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ and $\gamma$ denotes the discount factor. In an MDP, we aim to maximize the expected cumulative reward values, which are defined as $Q^{\pi}(s,a)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r\left(s_{t},a_{t}\right)\bigg{|}\pi,s_{0}=s,a_{0}=a\right]$ and $V^{\pi}(s)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r\left(s_{t},a_{t}\right)\bigg{|}\pi,s_{0}=s\right]$ for state-action pairs and states, respectively. Moreover, we have the advantage function $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$ Thus, the learning problem is formulated as Equation (1), where $\rho$ denotes an initial state distribution.
@@
In our framework, reward values are assigned by reviewer LLMs, which assess the diversity of actions within a trajectory. A trajectory with highly diverse actions receives a higher reward. The state $s_{t}\in\mathcal{S}$ on a trajectory $\mathcal{T}$ represents the generative instructions. Action space $\mathcal{A}$ is designed in Sec. 3.1.1, and specifical reward settings are provided in Sec. 3.1.2. On the basis of these settings, we can employ a model-free RL algorithm (e.g., TRPO Schulman et al. (2015)) to train the instructor LLM.
@@ 3.1.1 Action Set
Inspired by WizardLM (Xu et al., 2024), we leverage several actions to generate complex instructions, including “breadth action”, “add constraint”, “deepening”, “concretizing”, “increase reasoning steps”, “complicate input” (Xu et al., 2024). However, distinct from WizardLM, we map this discrete set into a continuous space by using a language encoding to represent each action, which is then used for a policy to generate instructions. This enables inherently capturing contextual nuances for direct comparison between actions via their Q values. The details of each action are in Appendix C.
@@ 3.1.2 Reward Settings
In this section, consider an evaluation prompt $g$. If $g$ is characterized as “equal”, the corresponding reward $r$ is designated as 0. Otherwise, $r$ is set to 1. The term “equal” here refers to semantic equivalence. It is noteworthy that the judgment is determined by the reviewer LLM. For instance, when conformity exists between the initial and subsequent instructions, the reward $r$ is 0. In contrast, a lack of conformity yields a reward value of 1.
For example, an evaluation prompt = “““ Here are two Instructions to ChatGPT AI, do you think they are equal to each other, which meet the following requirements: 1. They have same constraints and requirments. 2. They have same depth and breadth of the inquiry. The First Instruction: “““ +instruction + ”””. The Second Instruction: “““ + state + ”””. Your Judgement (Must Just only answer: Equal or Not Equal. No need to explain the reason; ‘Equal’ and ‘Not Equal’ are not allowed to appear simultaneously.): ”””.
@@ 3.2 Fine-tune a Pre-Aligned LLM
As illustrated in Fig. 1, an approach is adopted to optimize cost-efficiency considerations during the phase dedicated to fine-tuning a pre-aligned LLM. Specifically, we use the initial instructions from the Alpaca dataset.
The fine-tuning process is executed as follows: We first employ the Alpaca dataset [^4] https://huggingface.co/datasets/tatsu-lab/alpaca as the initial set of instructions and input the initial instructions into the expert LLM like ChatGPT. Next, we apply our trained instruction generation policy (derived from the Instructor LLM) to guide the Expert LLM (ChatGPT) in generating high-quality instructions. Specifically, this policy acts as a teacher directing the Expert LLM to produce instructions that meet our quality criteria. The instructor LLM provides different responses during the RL training process. Expert LLM is used to generate instructions and their associated responses, it serves as the source of knowledge and expertise. Then, the expert LLM produces responses to these high-quality instructions. With the instructions and responses in hand, we fine-tune the pre-aligned LLMs like Llama-1-chat-7b and Llama-2-chat-7b via a supervised fine-tuning way (SFT)—a cost-effective way to enhance LLM capabilities. Finally, we can have the post-SFT LLMs, such as TeaMs-RL-1-7b.
@@ 3.3 Practical Algorithms
The practical training process is outlined in Algorithm 1, which describes the training pipeline. A more detailed algorithm can be found in Algorithm 2, provided in Appendix B. In our training process, in order to teach expert LLMs how to generate high-quality instructions, we leverage the RL policy we have trained to prompt expert LLMs to generate instructions step by step. Building on the training process, we fine-tuned an LLM called “TeaMs-RL”. Note that we have the flexibility to choose any other RL methods for policy training, here we choose TRPO (Schulman et al., 2015) (TRPO is not the only choice), though it has longer training times versus alternatives like Proximal Policy Optimization. We select TRPO for its rigorous advantage function handling and theoretical guarantees of monotonic improvement.
@@ 4.1 Enhancement of Instruction Diversity
Figure 3: Comparison of our method with WizardLM in terms of data diversity.
We deploy a policy that is designed to orchestrate a trajectory consisting of six distinct instruction actions, in accordance with the training process to facilitate the generation of instructions by expert LLMs. Note, the expert LLMs utilized in our study, as well as those used by WizardLM, are ChatGPT-3.5 and ChatGPT-4. Both models were accessed in 2023, ensuring that we employed the same versions of the ChatGPT-3.5 and ChatGPT-4 models to generate instructions and responses. A noteworthy aspect of our investigation involves a comparative assessment of the data quality resulting from the utilization of our policy, in contrast to the approach adopted by WizardLM (Xu et al., 2024), which relies upon random sampling for data generation by querying expert LLMs. It bears emphasizing that our instruction actions are randomly initialized, but as training progresses, our policy iteratively learns to enable the instructor LLM to produce increasingly complex and diverse instructions.
As shown in Fig. 3, our method reliably enhances the diversity score of the instruction set. [^5] In designing instruction actions, the “breadth action” is a single action regarding breath thoughts. Thus, we insert the action into the middle of our trajectory to enhance breadth instructions after training a policy. The main computation overhead is learning the instruction policy on the relatively small WizardLM-13b in less than 1 hour on 2 NVIDIA RTX A6000 GPUs (with 896 total queries). This results in a transferable policy that reduces alignment cost across models compared to RLHF’s per-model RL. Offloading to policy learning provides an instruction set for joint tuning and alignment—a substantial benefit over tuning-only data usage. For detailed information on how to teach expert LLMs to generate instructions, please refer to Appendix D.
@@ 4.2 Compare with WizardLM-7b on ARC and HellaSwag Benchmakrs
To comprehensively examine the effectiveness of our method, we carry out experiments on popular benchmarks: (1) AI2 Reasoning Challenge (ARC) benchmark (Clark et al., 2018): The benchmark introduces a fresh question set and baselines, all strategically curated to foster AI research in the realm of advanced question answering, setting a significantly higher bar for knowledge and reasoning capabilities compared to previous challenges. (2) HellaSwag benchmark (Zellers et al., 2019): The benchmark introduces a challenging dataset, revealing that even state-of-the-art models struggle with commonsense inference, as evidenced by the significant performance gap between humans ($95\%$ accuracy) and models ($48\%$), achieved through adversarial filtering, a robust data collection paradigm that selects adversarial machine-generated wrong answers by scaling up the length and complexity of dataset examples to a ‘Goldilocks’ zone where the text generated is absurd to humans yet often misclassified by models.
We trained a llama-1-7b model denoted “TeaMs-RL-7b-v1.1” with our dataset of 17,878 instruction-response pairs, the training time is about 2 hours on 4 NVIDIA RTX A6000 GPUs. As shown in Figures 4 and 5, our method shows superior performance over WizardLM-7b models [^6] https://huggingface.co/TheBloke/wizardLM-7B-HF in the same experimental settings. It is crucial to highlight an important aspect of our methodology in relation to data utilization. The dataset employed for training our model is approximately one-fourteenth the size of the dataset utilized by WizardLM, as illustrated in Figure 5 (a). Furthermore, Fig. 5 (b) highlights the discernible difference in the query count posed to ChatGPT models between our method and WizardLM, with the latter requesting ChatGPT for responses more than seventeen times than TeaMs-RL. This marked contrast underscores the cost-effectiveness of our data generation approach, which mitigates the expenses associated with dataset acquisition and generates high-quality data. It highlights that our method is a more economically viable and sustainable strategy for training LLMs. For more experimental results, please see Appendix G. Note that the ChatGPT models we used to query data show performance equivalent to the ChatGPT models used in the baselines.
Figure 4: Comparison of our method with WizardLM 7B on LM-Eval Benchmarks (the higher the value, the better the method’s performance).
Figure 5: Comparison of our method with WizardLM 7B on dataset size used for training LLMs (a) and querying count of advanced LLMs (b) (the lower the value, the better the method’s performance).
@@
4.3 Model Privacy Attack Experiments